Find the whole quantity if 
  12% of it is ₹ 1080.
Find the whole quantity if 
  12% of it is ₹ 1080.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the whole quantity be x, Then,
(12/100) × (x) = 1080
X = 1080 × (100/12)
X = 540 × (100/6)
X = 90 × 100
X =₹9000

Thanks